PART A: The electrons in solids can be found ____________
in only certain discrete sharp energy states associated with their orbits.
in energy states that overlap so that more than one electron is associated with a given energy level.
in the same energy states as if the atoms forming the solid were far enough so that their interactions could be neglected.
in closely spaced energy levels that form a continuous distribution of energy within a certain range.


PART B: When an electron in the valence band in an insulator gains additional energy, it can jump to ___________
an adjacent energy state within the valence band.
an adjacent energy state in the conduction band.
an energy state in the conduction band only if it has gained an amount of energy at least equal to the band gap.
an energy state between the valence band and the conduction band.

PART C: When an electron in the highest energy band in a conductor gains additional energy, it can jump to ___________
an adjacent energy state within the same band.
an energy state in the next-higher band.
an energy state just above the conduction band.

PART D: What is the main difference between insulators and semiconductors?
In semiconductors, the valence band is only partially filled.
In semiconductors, the energy gap between the valence and conduction band is considerably smaller than in insulators.
At absolute zero, in semiconductors some electrons occupy energy states in the conduction band.

PART E: The energy gap between the valence band and the conduction band in germanium is only 0.67 . If a germanium crystal is irradiated with electromagnetic waves of 1.37×10-6 wavelength, will some electrons in the top level of the valence band gain enough energy to jump to the lowest energy level in the conduction band?
yes
no
Answer #1

PART A:

The electrons in solids can be found

"in closely spaced energy levels that form a continuos distribution of energy with in a certain range."

PART B:

When an electron in the valance band in an insulator gains additional energy.

it can jump to

"an energy state in the conduction band only if it has gained an amount of energy at least equal to the band gap."

PART C:

When an electron in the highest energy band in a conductor gains additional energy, it can jump to

an adjacent energy state within the same band.

PART D:

The main difference between insulators and semiconductors is that

"In semiconductors, the energy gap between the valence and conduction

band is considerably smaller than in insulators."

PART E:

The energy arsociated with this wave length is

$$ \begin{aligned} &\Delta \mathrm{E}=\frac{h c}{\lambda}=\frac{\left(6.63 \times 10^{34} \mathrm{Js}\right)\left(3 \times 10^{6} \mathrm{~m} / \mathrm{s}\right)}{\left(1.37 \times 10^{-6} \mathrm{~m}\right)} \\ &\Rightarrow \Delta \mathrm{E}=14.518 \times 10^{30} \mathrm{~J} \end{aligned} $$

Converting this in \(\mathrm{eV}\), we have \(\Delta \mathrm{E}=\frac{14.518 \times 10^{.20}}{1.6 \times 10^{-8}}\)

\(\Rightarrow \Delta \mathrm{E}=0.907 \mathrm{eV}>\) the band gap of germanium \(=0.67 \mathrm{eV}\)

Hence the elecrons will jump from valance band to conduction band

Hence the answer is YES.